What is the molarity of Fe(CN)6^4- is saturated solution of Ag4(Fe(CN)6)?ksp = 1.6*10^-41

Grade:

1 Answers

Vikas TU
4 years ago
Dear student 
 
According to the theory:-
{Ag}{Fe(CN)6}=Ksp
x*4^4*x^4=Ksp
256x^5=1.6*10^-41
x^5=6.25*10^-44
Therefore x=2.3*10^-9
Hope this helps .
Good luck 
Cheers